OpenStudy (anonymous):

What is a series like below called? I don't know how to approach it... (Apologies for too many parentheses and any typos, I'm not sure how to write this clearly in this website) 1st term: 1/(1+(1/(2+1))) 2nd term: 1/(1+(1/(2+(1/(1+(1/(2+1)))))) 3rd term: 1/(1+(1/(2+(1/(1+(1/(2+(1/(1+1/(2+1)))))))))) In each successive term the last "1" turns into another 1/(1+1/(2+1))

OpenStudy (anonymous):

continued fraction?

OpenStudy (anonymous):

Is there a way to see if it approaches a particular limit as you add successive terms?

OpenStudy (anonymous):

\[\frac{1}{1+\frac{1}{3}}=\frac{1}{\frac{4}{3}}=\frac{3}{4}\] is that the first one? it is hard for me to read

OpenStudy (anonymous):

Yes, that's the first one. The next one ends up 11/15, then 41/56, then 153/209, if my napkin-math is right.

OpenStudy (anonymous):

gimmick is to call the limit \(x\) and then see if you can write an expression in \(x\)

OpenStudy (anonymous):

Mm, I played around with that some, but didn't see how to do it.

OpenStudy (phi):

\[ x= \frac{1}{1+\frac{1}{2+x}} \]

OpenStudy (anonymous):

Ah, that looks simple. Thank you both!

OpenStudy (anonymous):

Hrm... is there a way to give you both credit for helping? I only see a "best response".

OpenStudy (phi):

I hope it makes sense... x= continued fraction forever. replace the part of it with the continued fraction..

OpenStudy (phi):

no. But sat needs the medals.

OpenStudy (anonymous):

try \[x=\frac{1}{1+\frac{1}{2+x}}\] but maybe this is wrong, not clear

OpenStudy (anonymous):

It does make sense - so I can just look at how the fraction behaves with really large values of x. And I'll be happy to give the medal to Sat, then.

OpenStudy (anonymous):

oops what @phi said yeah, i need the medals, going to buy my car with them!

OpenStudy (phi):

@sat yes, that works....once you solve you can check against 153/209

OpenStudy (anonymous):

I'm playing around with an EE question, and wandered into this. Thank you much!

OpenStudy (phi):

you solve for x x= -1±sqrt(3)

OpenStudy (anonymous):

That looks right. Awesome.